OpenStudy (anonymous):

Let f be the function given by square root of x^4-16x^2 a) Find the domain of f. b) Find f '(x) c) Find the slope of the line normal to the graph of f at x = 5

OpenStudy (rulnick):

a. all real numbers b. 4x^3-32x c. f'(5)=4(5)^3-32(5)=500-160=340, so normal slope is -1/340

OpenStudy (anonymous):

thank you :)

OpenStudy (rulnick):

welcome

OpenStudy (anonymous):

7th time where is this question coming from?

OpenStudy (anonymous):

domain is not all real number becauase it is square root of something, not "something"

OpenStudy (anonymous):

but this is not the first time i have seen this exact same question not the second time either. would like to know where it comes from, and i can give a complete answer

OpenStudy (anonymous):

@tserio94 please let me know where this question comes from

OpenStudy (anonymous):

what do you mean where does it come from?

OpenStudy (anonymous):

i mean what text, homework assignment etc? where does it come from? the answer is not all real numbers for the domain

OpenStudy (anonymous):

yeah its a hw assignment

OpenStudy (anonymous):

from what text?

OpenStudy (anonymous):

or did your math teacher assign it?

OpenStudy (anonymous):

she just assigned it

OpenStudy (anonymous):

did she ask you to post it on open study? this is the 8th time i have seen this same question

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that is what i thought

OpenStudy (anonymous):

do you know why she asked you to do that?

OpenStudy (anonymous):

an experiment maybe?

OpenStudy (anonymous):

i dont know why to be honest

OpenStudy (anonymous):

is it a hight school calculus class or college?

OpenStudy (anonymous):

highschool

OpenStudy (anonymous):

cool are you taking calc?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

nice what this the questinon \[\sqrt{x^4-16x^2}\]

OpenStudy (anonymous):

to find the domain of this function?

OpenStudy (anonymous):

i cant see it says [math processing error]

OpenStudy (anonymous):

yeah it is not working so well tonight square root of x^4-16x^2 maybe?

OpenStudy (anonymous):

yeah how would you find that

OpenStudy (anonymous):

you would have to make sure that x^4-16x^2 was greater than or equal to zero, can you do that?

OpenStudy (anonymous):

i think , would you have to distibute the x^4-16x^2 right?

OpenStudy (anonymous):

no the first step would be to factor

OpenStudy (anonymous):

can you factor this one?

OpenStudy (anonymous):

my bad thats what i meant

OpenStudy (anonymous):

and i got x^2(x^2-16)

OpenStudy (anonymous):

would it be now x^2(x-4)(x+4) ?

OpenStudy (anonymous):

yes that would work

OpenStudy (anonymous):

now we know that x^2 is always greater than or equal to zero right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so that leaves (x-4)(x+4) do you know for which values of x this is positive?

OpenStudy (anonymous):

i got -4 and 4

OpenStudy (anonymous):

that is wehre it is zero, but you need to know where it is positive

OpenStudy (rulnick):

Ah, okay, stepped out. Surprised to see the long discussion. I missed the "sqrt of" part. Sorry. Thanks s73 for catching this.

OpenStudy (anonymous):

no problem, i am happy to help just wondering what school this is because this is a pretty sophisticated problem for high school

OpenStudy (anonymous):

its an ap class and idk what you mean where is it possitive

OpenStudy (anonymous):

positive means greater than zero. you cannot take the square root of a negative number so you have to make sure the expression inside the radical is not negative

OpenStudy (anonymous):

ap calc sounds hard

OpenStudy (rulnick):

a. all real x with |x| > 4 (x less than -4 or greater than 4) b. (2x^3-16x)/sqrt(x^4-16x^2) c. f'(5) = 34/3, normal slope is -3/34

OpenStudy (anonymous):

now one factor is x^2so you don't have to worry about it, becuase it is never negative, but (x-4)(x+4) could be negative

OpenStudy (rulnick):

a. detail correction: should be >= (so x <= -4 or x >= 4)

OpenStudy (anonymous):

wait so x|-4<x>4 ?

OpenStudy (anonymous):

no

OpenStudy (rulnick):

When I gave you (a) above I said: a. all real x with |x| > 4 (x less than -4 or greater than 4) but I should have said a. all real x with |x| >= 4 (x less than or equal to -4 or greater than or equal to 4)

OpenStudy (anonymous):

x has to be less than -4 or greater than 4 it could also be 0

OpenStudy (rulnick):

yes, good point (no pun intended), x=0 is in the domain as well

OpenStudy (anonymous):

but how do i write that as an answer?

OpenStudy (rulnick):

domain = {x in R | x=0 or |x|>=4}

OpenStudy (rulnick):

... people here seem to like "|" but ":" is better to avoid confusion, so: domain = {x in R : x=0 or |x|>=4}

OpenStudy (anonymous):

but what about the -4 what happened to that?

OpenStudy (rulnick):

It's in there. If |x| >= 4 then x>=4 or x<=-4.

OpenStudy (anonymous):

if |x|>4 that is a succinct way to write x<-4 or x > 4

OpenStudy (rulnick):

They're equivalent.

OpenStudy (anonymous):

what notation do they use in ap calc?

OpenStudy (anonymous):

so its x|x=o or x>4 or x< -4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you know how to find the derivative?

OpenStudy (anonymous):

but isnt there a shprter neater way to write that or it is good as is? and yes , but since theres a sq root wed have to use chain rule right?

OpenStudy (anonymous):

you can write it in simplest radical form first

OpenStudy (anonymous):

but it is probably easiest to use the chain rule

OpenStudy (anonymous):

ok so would it be 2x^3-16x/ sq root x^4-16^2

OpenStudy (rulnick):

There is one nice simplification you can use, if you know this: sgn(x) = x / |x| It's the "signum" function and returns the sign of its argument: sgn(x) = 0 if x=0 = 1 if x>0 = -1 if x<0 This would allow you to reduce the derivative to 2sgn(x) (x^2-8) / (x^2-16)^(1/2)

OpenStudy (anonymous):

yeah and you can divide by 2 and also divide by x, by you have to be careful as rulnik said, because you don't know if x is positive or negative

OpenStudy (anonymous):

so you if you simplify this way you have to take in to accound whether x is positive or negative

OpenStudy (anonymous):

is it ok to just leave it w/the answer i got?

OpenStudy (anonymous):

that is a good idea

OpenStudy (rulnick):

I agree.

OpenStudy (anonymous):

dont forget that the derivative of the square root of something should have a two in the denominator, so you should divide the numerator by 2 to get 2x^2 - 16x

OpenStudy (anonymous):

sorry i mean 2x^3-16x

OpenStudy (anonymous):

i am still wondering where such a difficult question came from

OpenStudy (anonymous):

from a book or from a math teacher

OpenStudy (rulnick):

I think that's what tserio had above, (2x^3-16x)/sqrt(x^4-16x^2) [a typo or two fixed here], so should be good to go I think.

OpenStudy (rulnick):

I'd be curious to know if tserio's teacher catches the details as well as you two did :) Nice work.

OpenStudy (anonymous):

i would be interested to know who it is

OpenStudy (anonymous):

is part b good? how about the last part?

OpenStudy (anonymous):

did your math teacher give you any other questions to ask here?

OpenStudy (anonymous):

yes but what would be the steps needed to solve part c of this question

OpenStudy (anonymous):

you would have to replace x by 5 to see what you get what other questions?

OpenStudy (anonymous):

is this an on line class or a regular one?

OpenStudy (anonymous):

plug 5 into sq root of x^4-16x^2? and no regular class

OpenStudy (rulnick):

Yes, (2x^3-16x)/sqrt(x^4-16x^2) with x=5 is 34/3.

OpenStudy (anonymous):

ok so for normal you basically plug in and do reciprical?

OpenStudy (rulnick):

Yes.

OpenStudy (rulnick):

Negative reciprocal.

OpenStudy (anonymous):

ok thank you do you think you can help me with the other question i posted?

OpenStudy (rulnick):

where?

OpenStudy (anonymous):

Let R be the region in the xy-plane between the graphs of y = e^x and y = e^-x from x = 0 to x = 2. a) Find the volume of the solid generated when R is revolved about the x-axis.

OpenStudy (rulnick):

Will you join me at the other problem?